# cuTeSpMM: Accelerating Sparse-Dense Matrix Multiplication using GPU Tensor Cores

### v1

@@ 4. Analysis and Parameter Selection
The cuTeSpMM algorithm has parameters, (${TM}$, ${TK}$) representing the tile sizes along the dimensions $M$ and $K$ of blocks within a row panel. The $TK$ size will not have an affect on $\alpha$, thus, it doesn’t directly impact data movement and the number of FMA computations. However, it does influence other performance factors. In line 25 of our kernel design algorithm, we traverse all ${\frac{T_{K}}{brick_{k}}}$ brick columns inside a block. This for loop is fully unrolled since both ${TK}$ and ${brick_{k}}$ are constants. The prefix scanning operations used to decode the non-zero distribution pattern are entirely independent. A larger ${TK}$ value enhances ILP (Instruction-Level Parallelism), potentially improving performance. However, a larger ${TK}$ value also increases the volume of shared memory required, and thus reduces occupancy(line 3 in algorithm 1).Increased shared memory usage can reduce occupancy, potentially degrading parallelism. Quantifying ILP is challenging as the compiler affects the generated code. We empirically determined that a ${TK}$ size of 16 balances the occupancy and ILP.
In the previous ${OI}$ analysis, we assumed that ${TM}$ equals ${brick_{m}}$, signifying that every time we load a brick of the sparse A matrix, we must load the corresponding elements of B from shared memory to register in order to perform tensor core HMMA. If ${TM}$ is larger than ${brick_{m}}$, each brick column may contain ${\frac{TM}{brick_{m}}}$ bricks. The active bricks in a brick column require the same elements from the B matrix. Consequently, the B elements loaded from shared memory to register can be reused across all active bricks within a brick column. Let’s denote $\beta$ as the average number of bricks inside a brick column of a block. The number of shared memory transfers of the B matrix can then be reduced to:
@@
assuming the brick column density $\alpha$ remains constant. Thus, increasing $T_{M}$ directly increases the $OI_{shmem}$ due to increased reuse. However, this increase in reuse is often counteracted by a decrease in the column non-zero density $\alpha$ as $T_{M}$ increases. For a sparse matrix, a large ${TM}$ often leads to a drop in brick column density. Figure 8 illustrates two row panels, each with a height of ${TM}$ but featuring different non-zero distributions. The dashed line divides a row panel into two sub-row panels, each with a height of ${\frac{TM}{2}}$. In the first row panel, as depicted in Figure 8 (a), the active columns of the two sub-row panels are common. Left active column packing in this scenario results in a densely packed region highlighted in light blue. However, in the second row panel (Figure 8 (b)), the active columns of the two sub-row panels are disjoint. Consequently, after left packing, the dense region for the second row panel is much sparser than that of the first row panel. Furthermore, since the active columns of the two sub-row panels are entirely disjoint, there is no data reuse across the two sub-row panels. Hence, for the second row panel, it is more beneficial to set the row panel size to ${\frac{TM}{2}}$ and assign the two sub-row panels to different thread blocks. These two row panels represent the two extreme cases of different non-zero distributions within a row panel. In reality, the non-zero distribution of a row panel typically falls somewhere between these two extreme cases. Although a large ${TM}$ value has the potential to increase the data reuse of ${B}$ elements, thereby increasing $OI$, it also comes with a disadvantage. This disadvantage arises from the fact that a large ${TM}$ increases shared memory and register usage, which degrades occupancy. Consider that the matrices from suitesparse are super sparse, and the non-zero distribution is far away from the distribution shown in Figure 8 a, and the disadvantage of the occupancy drop for large $TM$. In our evaluation, all the matrices we evaluated are using $TM$ = 16, which is the size of ${brick_{m}}$.
@@ 5. Load balancing improvement
Inspired by this example, we design our load balancing scheme as follows: First, for each row panel, we compute the $num\_loads$ of that row panel. The $num\_loads$ is defined as the equation:
@@
Then, for each row panel, we partition it into multiple row panels based on the $partition\_ratio$ (no partition happens if $partition\_ratio=1$).
To compute the $num\_waves$ in equation 3, we need to determine the number of thread blocks that can be resident in an SM concurrently and the number of SMs. This information can be obtained during the compilation phase by querying the device (GPU) information. We also need the ${Total\_thread\_blocks}$ which is determined at runtime.
Compared with the second approach, which partitions the row panel only based on $average\_blocks\_each\_row\_panel$, our strategy reduces the partition ratio by the factor of $num\_waves$, translating to a reduction in the number of atomic operations by the factor of $num\_waves$.
@@ 6.4. Sparse Matrix Synergy
Table 2 shows the distribution of the tested Suitesparse matrices in terms of these three degrees of synergy for TCU-based SpMM execution.
Matrix n=32 n=64 n=128
cuTeSpMM
TC-GNN
Best-SC
cuTeSpMM
TC-GNN
Best-SC
cuTeSpMM
TC-GNN
Best-SC
amazon0505 857 122 811 1024 202 746 1077 271 940
amazon0601 815 161 991 988 274 669 1049 383 899
artist 510 73 626 655 135 864 782 202 1361
citeseer 56 13 41 105 26 93 176 47 152
com-amazon 455 98 467 587 170 447 657 249 616
cora 59 10 45 113 20 81 205 37 150
DD 821 177 767 1034 302 560 1150 426 925
OVCAR-8H 497 117 396 611 196 454 674 275 508
ppi 587 107 720 717 198 708 929 297 1235
PROTEINS_full 457 126 515 690 220 445 788 330 633
so pubmed 269 43 233 359 79 325 502 131 581
soc-BlogCatalog 863 69 638 1079 126 989 1374 191 1644
Yeast 507 125 401 627 209 454 695 295 507
YeastH 499 118 393 615 198 457 667 276 495
Table 3. Comparing the performance (GFLOPs) of cuTeSpMM, TC-GNN, and Best-Scalar for matrices tested in (wang2023tc, )
Matrix n=32 n=128 n=512
cuTeSpMM
Best-SC
TC-GNN
cuTeSpMM
Best-SC
TC-GNN
cuTeSpMM
Best-SC
TC-GNN
amazon0505 848 811 122 1,077 940 271 1,088 957 900
amazon0601 807 991 161 1,047 899 383 1,059 920 1,172
artist 512 626 73 756 1,361 202 929 1,532 680
citeseer 53 41 13 157 152 47 315 292 156
com-amazon 446 467 98 653 616 249 689 620 767
cora 55 45 10 206 150 37 443 356 125
DD 808 767 177 1,143 925 426 1,171 985 1,206
OVCAR-8H 495 396 117 671 508 275 667 496 696
PROTEINS_full 447 515 126 822 633 330 963 789 912
so pubmed 272 233 43 528 581 131 651 770 419
soc-BlogCatalog 828 638 69 1,360 1,644 191 1,301 1,882 662
Yeast 526 401 125 692 507 295 685 493 740
YeastH 507 393 118 664 495 276 664 482 671
Table 4. Comparing the Ampere A100 performance (GFLOPs) of cuTeSpMM, TC-GNN, and Best-Scalar for matrices evaluated in TCGNN paper
@@ 7.3. FLASH LLM
Recently, Xia et al. introduced $FLASH\_LLM$ SpMM in their paper (xia2023flash, ), leveraging Tensor Cores to accelerate the SpMM operation within the machine learning domain. Notably, the matrices they evaluated were self-generated random matrices with densities ranging from 10% to 30%, which are considerably denser than the matrices typically encountered in GNN/scientific computing, our primary focus. Additionally, they employed a different data type, FP16, while we utilize FP32 (TF32). The matrices we target tend to have densities below 1%, presenting further challenges such as low computation efficiency (low brick density) and load imbalance (non-zeros are not uniformly distributed), which must be addressed to achieve optimal performance. We were unable to directly compare performance with their implementation due to the difference in data precision targeted. Further, the domain they target features sparse matrices with very much lower sparsity, typically between 60% to at most 95%.

### v2

@@ 4. Analysis and Parameter Selection
The cuTeSpMM algorithm has parameters, (TM{TM}, TK{TK}) representing the tile sizes along the dimensions $M$ and $K$ of blocks within a row panel. The $TK$ size will not have an affect on $\alpha$, thus, it doesn’t directly impact data movement and the number of FMA computations. However, it does influence other performance factors. In line 25 of our kernel design algorithm, we traverse all $\frac{T_{K}}{brick_{k}}$} brick columns inside a block. This for loop is fully unrolled since both TK{TK} and $brick_{k}$} are constants. The prefix scanning operations used to decode the non-zero distribution pattern are entirely independent. A larger TK{TK} value enhances ILP (Instruction-Level Parallelism), potentially improving performance. However, a larger TK{TK} value also increases the volume of shared memory required, and thus reduces occupancy(line 3 in algorithm 1).Increased shared memory usage can reduce occupancy, potentially degrading parallelism. Quantifying ILP is challenging as the compiler affects the generated code. We empirically determined that a TK{TK} size of 16 balances the occupancy and ILP.
In the previous OI{OI} analysis, we assumed that TM{TM} equals $brick_{m}$}, signifying that every time we load a brick of the sparse A matrix, we must load the corresponding elements of B from shared memory to register in order to perform tensor core HMMA. If TM{TM} is larger than $brick_{m}$}, each brick column may contain $\frac{TM}{brick_{m}}$} bricks. The active bricks in a brick column require the same elements from the B matrix. Consequently, the B elements loaded from shared memory to register can be reused across all active bricks within a brick column. Let’s denote $\beta$ as the average number of bricks inside a brick column of a block. The number of shared memory transfers of the B matrix can then be reduced to:
@@
assuming the brick column density $\alpha$ remains constant. Thus, increasing $T_{M}$ directly increases the $OI_{shmem}$ due to increased reuse. However, this increase in reuse is often counteracted by a decrease in the column non-zero density $\alpha$ as $T_{M}$ increases. For a sparse matrix, a large TM{TM} often leads to a drop in brick column density. Figure 8 illustrates two row panels, each with a height of TM{TM} but featuring different non-zero distributions. The dashed line divides a row panel into two sub-row panels, each with a height of $\frac{TM}{2}$}. In the first row panel, as depicted in Figure 8 (a), the active columns of the two sub-row panels are common. Left active column packing in this scenario results in a densely packed region highlighted in light blue. However, in the second row panel (Figure 8 (b)), the active columns of the two sub-row panels are disjoint. Consequently, after left packing, the dense region for the second row panel is much sparser than that of the first row panel. Furthermore, since the active columns of the two sub-row panels are entirely disjoint, there is no data reuse across the two sub-row panels. Hence, for the second row panel, it is more beneficial to set the row panel size to $\frac{TM}{2}$} and assign the two sub-row panels to different thread blocks. These two row panels represent the two extreme cases of different non-zero distributions within a row panel. In reality, the non-zero distribution of a row panel typically falls somewhere between these two extreme cases. Although a large TM{TM} value has the potential to increase the data reuse of B{B} elements, thereby increasing $OI$, it also comes with a disadvantage. This disadvantage arises from the fact that a large TM{TM} increases shared memory and register usage, which degrades occupancy. Consider that the matrices from suitesparse are super sparse, and the non-zero distribution is far away from the distribution shown in Figure 8 a, and the disadvantage of the occupancy drop for large $TM$. In our evaluation, all the matrices we evaluated are using $TM$ = 16, which is the size of $brick_{m}$}.
@@ 5. Load balancing improvement
Inspired by this example, we design our load balancing scheme as follows: First, for each row panel, we compute the num_loadsnum\_loads of that row panel. The num_loadsnum\_loads is defined as the equation:
@@
Then, for each row panel, we partition it into multiple row panels based on the partition_ratiopartition\_ratio (no partition happens if partition_ratio=1partition\_ratio=1).
To compute the num_wavesnum\_waves in equation 3, we need to determine the number of thread blocks that can be resident in an SM concurrently and the number of SMs. This information can be obtained during the compilation phase by querying the device (GPU) information. We also need the Total_thread_blocks{Total\_thread\_blocks} which is determined at runtime.
Compared with the second approach, which partitions the row panel only based on average_blocks_each_row_panelaverage\_blocks\_each\_row\_panel, our strategy reduces the partition ratio by the factor of num_wavesnum\_waves, translating to a reduction in the number of atomic operations by the factor of num_wavesnum\_waves.
@@ 6.4. Sparse Matrix Synergy
Table 2 summarizes the definition of the three synergy levels, while table 2 shows the distribution of the tested Suitesparse matrices in terms of these three degrees of synergy for TCU-based SpMM execution.
@@ 7.3. FLASH LLM
Recently, Xia et al. introduced FLASH_LLMFLASH\_LLM SpMM in their paper (xia2023flash, ), leveraging Tensor Cores to accelerate the SpMM operation within the machine learning domain. Notably, the matrices they evaluated were self-generated random matrices with densities ranging from 10% to 30%, which are considerably denser than the matrices typically encountered in GNN/scientific computing, our primary focus. Additionally, they employed a different data type, FP16, while we utilize FP32 (TF32). The matrices we target tend to have densities below 1%, presenting further challenges such as low computation efficiency (low brick density) and load imbalance (non-zeros are not uniformly distributed), which must be addressed to achieve optimal performance. We were unable to directly compare performance with their implementation due to the difference in data precision targeted. Further, the domain they target features sparse matrices with very much lower sparsity, typically between 60% to at most 95%.
@@ 8. Conclusion
Acknowledgements.
This work was supported in part by the U.S. National Science Foundation through award 2009007.